### v1

@@ 4.2.5 Speech-Based Personalization
4.2.6 Robustness in Noisy Environments
@@
4.3 Educational and Cultural Studies
@@
4.3.1 Language Learning Tools
@@
4.3.2 Children’s Language Development
@@
4.3.3 Cultural Analysis and Documentation
@@
4.3.4 Applications in Linguistic and Cultural Heritage Projects
@@
5 Challenges and Limitations
@@
5.1 Data Scarcity and Representativeness
@@
  • Restricted Access. A significant number of existing datasets are only available for a fee or under restrictive licensing agreements, limiting their accessibility to researchers and developers with substantial funding. For example, commercial datasets such as Aurora Project [28] Reference] and Defined.ai [7, 3, 25] Italian Datasets require significant financial investment, making them inaccessible for smaller research teams or open-source projects. This restricts innovation and reinforces inequalities in data access. In Table 2, we present a list of datasets that are available under different forms of “open access” together with some relevant features.
@@
Table 2: Summary of Publicly Available Italian Speech Datasets
@@ 5.3 Ethical Considerations
6 Future Directions and Recommendations
@@
7 Conclusion
Spoken Italian datasets have become indispensable tools across diverse domains, from linguistic research to cutting-edge speech technologies. Their ability to capture the rich diversity of Italian speech—spanning regional dialects, sociolinguistic variations, and diverse speech contexts—enables researchers and developers to address complex linguistic questions and build robust applications for real-world use.
This survey highlights the critical contributions of these datasets while addressing the challenges that limit their current potential. Issues such as data scarcity, accessibility, and ethical considerations underscore the need for collaborative and innovative approaches to dataset creation and dissemination. By prioritizing inclusivity, standardization, and open access, the research community can ensure that these resources remain relevant and impactful.
Moreover, advancements in data collection and annotation technologies present exciting opportunities for expanding the scope and quality of spoken Italian datasets. The integration of automated tools, multimodal data, and interdisciplinary collaboration will not only enhance dataset utility but also drive innovation in natural language processing, education, and cultural preservation. As Italy’s linguistic landscape continues to evolve, these datasets will play a vital role in documenting, understanding, and preserving its unique linguistic heritage. By addressing current gaps and exploring future possibilities, the development and application of spoken Italian datasets can contribute significantly to the advancement of both academic research and practical technologies, fostering a deeper appreciation of Italy’s linguistic and cultural richness.
@@ 7 Conclusion
To support the exploration and development of spoken Italian datasets, this survey highlights a curated collection of 66 datasets, encompassing diverse speech types, sources, and linguistic features. While this paper discusses key examples to illustrate the challenges, methodologies, and applications in the field, the full collection is made publicly available for further reference and research. The complete dataset repository, hosted on GitHub and archived on Zenodo (DOI: 10.5281/zenodo.14246196), aims to facilitate collaborative efforts and inspire future advancements in Italian linguistic research and speech technology.
Biographies
Marco Giordano is a Full Professor in Musical Acoustics and Perception at the Conservatory of Music of L’Aquila, Italy and currently PhD student in ICT and Telecommunications Engineering at the DISIM dpt. of the University of L’Aquila. He has a Laurea degree in Electronic Engineering from the University of Roma. His main research activities are focused on digital signal processing, musical acoustics and perception, digital education and deep neural models applied to computational paralinguistics.
Claudia Rinaldi is a Researcher in Signal Processing at CNIT (National Inter-University Consortium for Telecommunications) and Adjunct Professor at the Department of Information Engineering, Computer Science and Mathematics of the University of L’Aquila, Italy. She received the Laurea degree (cum laude) and Ph.D. degree in Electronic Engineering from the University of L’Aquila, Italy, in 2005 and 2009, respectively. She also got a master degree in Trumpet and bachelor degree in Electronic Music at the Conservatory of Music of L’Aquila in 2006 and 2013 respectively.
Her main research activities are focused on digital signal processing algorithms and more in general on the use of technology in artistic fields. Moreover, she is concerned with design, modeling and optimization of communication algorithms with particular emphasis on the physical layer and software defined radio for the development of transmission systems responding to cognitive radios paradigms.

### v2

@@ IV-B5 Speech-Based Personalization
IV-B6 Robustness in Noisy Environments
@@
IV-C Educational and Cultural Studies
@@
IV-C1 Language Learning Tools
@@
IV-C2 Children’s Language Development
@@
IV-C3 Cultural Analysis and Documentation
@@
IV-C4 Applications in Linguistic and Cultural Heritage Projects
@@
TABLE II: Datasets vs. Applications (Section IV). A check mark (✓) indicates relevance to that specific sub-subsection. The table refers to a subset (23 over 66) of reported datasets. The complete datasets collection can be found on GitHub.
4.1 Linguistic Research 4.2 NLP & Speech Tech 4.3 Educational & Cultural
Dataset Phonetics Phonology Syntax Morphology Sociol. Dialect. Discourse Pragmatics ASR TTS Machine Translation Sentim. Emotion Person. Noise Robustness Language Learning Tools Child.’s Language Dev. Cultural Analysis Heritage Projects
Apasci [41]
Aurora Project [28]
C-ORAL-ROM [14]
Common Voice [19]
DEMoS [33, 32]
DIA [4]
EMOFILM [30]
Europarl-ST [36, 37]
IBNC [24]
Italian Spontaneous Dialogue (Telephony) [27]
Italian Kids Speech Recognition Corpus [42]
Italian Lifestyle Podcast [25]
Italian Scripted Monologue [7]
Italian Spontaneous Dialogue Dataset [3]
KIParla [1, 2]
LABLITA [13, 12]
M-AILABS [39]
MLS [17, 18]
MULTEXT Prosodic [40]
MuST-C [29]
PortMedia [38]
VIVALDI [31]
VoxPopuli [8, 9]
V Challenges and Limitations
@@
V-A Data Scarcity and Representativeness
@@
  • Restricted Access. A significant number of existing datasets are only available for a fee or under restrictive licensing agreements, limiting their accessibility to researchers and developers with substantial funding. For example, commercial datasets such as Aurora Project [28] Reference] and Defined.ai [7, 3, 25] Italian Datasets require significant financial investment, making them inaccessible for smaller research teams or open-source projects. This restricts innovation and reinforces inequalities in data access. In Table III, we present a list of datasets that are available under different forms of “open access” together with some relevant features.
@@
TABLE III: Summary of Publicly Available Italian Speech Datasets
@@ V-C Ethical Considerations
VI Future Directions and Recommendations
@@
VII Conclusion
This work has presented a comprehensive survey of 66 spoken Italian datasets, examining their types, collection and annotation methodologies, as well as current limitations. In the introductory section (Section 1), we highlighted the gap between the resources available for Italian and those for major world languages such as English or Mandarin. Subsequently, we have proposed a classification of the datasets according to speech type (spontaneous, conversational, read, monologic), source and context (broadcast media, telephone, etc.), and demographic or linguistic features (dialects, sociolinguistic variation, multilingual settings).
In the successive Section, we focused on data collection and annotation strategies, underscoring the importance of standardized protocols, automated tools, and validation procedures (e.g., inter-annotator agreement) to ensure high-quality data. Section 4 has discussed key applications, ranging from automatic speech recognition (ASR) and text-to-speech (TTS) to emotion analysis and language learning, as well as cultural and heritage preservation.
We then addressed the main challenges and limitations: limited coverage of dialects and minority languages, restricted access and licensing issues, demographic imbalances, and ethical considerations related to privacy. Finally, Section 6 outlined several future recommendations, including a call for standardized annotation formats, closer collaboration between academic and industrial stakeholders, and the need for crowdsourced initiatives to broaden dialectal and linguistic coverage.
Based on these findings, the important actions to be carried on have been presented, such as expanding datasets to include underrepresented dialects and regional varieties, potentially by partnering with local communities for field recordings; promoting open access and revise data-sharing policies to enhance research and innovation, particularly for educational contexts and smaller enterprises; developing and adopting consistent annotation and validation standards (covering demographic, sociolinguistic, and prosodic metadata) to improve dataset interoperability; fostering interdisciplinary research and partnerships with industrial, cultural, and international academic institutions, aiming to advance Italian speech technologies on a global scale.
@@ VII Conclusion
To support the exploration and development of spoken Italian datasets, this survey highlights a curated set of 66 resources, encompassing diverse speech types, sources, and linguistic features. While this paper discusses selected examples to illustrate the challenges, methodologies, and applications in the field, the full collection is made publicly available for further reference and research. The complete dataset repository, hosted on GitHub and archived on Zenodo (DOI: 10.5281/zenodo.14246196), aims to facilitate collaborative efforts and inspire future advancements in Italian linguistic research and speech technology.